OpenStudy (anonymous):

simplify (x^2-16)/(4-x)

OpenStudy (curry):

(x+4)(x-4)/-(x+4) x-4/-1

OpenStudy (anonymous):

= (x-4)(x+4)/(4-x) = (x-4)(x+4)/-(x-4) = -(x+4) = -x-4

OpenStudy (curry):

wait is the negative sing allowed to be on the top/ cause i think it should stay at the denominator